# The Impact of Spectroscopic Redshift Errors on Cosmological Measurements

### v1

@@ 5.2.2 Catastrophics-correction for slitless survey
In this test, we present Full-Modeling fitting results based on the mean power spectrum calculated from all 500 mocks, with the covariance matrix rescaled to volume $V_{25}$. By averaging over 500 mocks, we minimize the impact of cosmic variance, so any observed deviations in parameter estimates can be more confidently attributed to redshift errors. The left panel of Figure˜6 shows the posterior distributions of cosmological parameters using the baseline EFT fit for three contamination cases: clean (no contaminations), ELG-like catastrophics ($f_{c}=1\%$), and slitless-like errors (i.e., $\sigma_{v}=85.7\rm{km}\cdot s^{-1}$ and $f_{c}=5\%$). The corresponding mean values with 68% confidence intervals for $\ln(10^{10}A_{s})$ are summarized in Table˜2. Compared the $\ln(10^{10}A_{s})$ constraint from clean case, we find a small shift of $0.36\sigma$ by ELG-like catastrophics and a significant deviation of $2.20\sigma$ by slitless-like errors. These results are consistent with our earlier findings in Section˜5.2.1.
To incorporate the effect of catastrophics in the clustering model, we extend the baseline EFT model by adding an additional parameter $f_{c}$ and multiplying the factor $(1-f_{c})^{2}$ to the galaxy power spectrum as described in Eq.˜4.6. We refer to the $\texttt{EFT}+{\rm{free}}\,f_{c}$ model, when the parameter $f_{c}$ varies during sampling with a flat prior $f_{c}\in[0,1]$. The posterior distributions, including the variation of $f_{c}$, are shown in the right panel of Figure˜6 (orange), with the corresponding constraint on $\ln(10^{10}A_{s})$ reported in Table˜2. The $\texttt{EFT}+{\rm{free}}\,f_{c}$ model is able to recover an unbiased mean value of $\ln(10^{10}A_{s})$ for slitless-like errors, with a discrepancy below $0.1\sigma$ compared to the clean case. However, this improvement comes at the cost of loss in constraining power: we measure a $\sim 60\%$ degradation in the $\ln(10^{10}A_{s})$ constraint. This loss is due to the strong degeneracy between $f_{c}$ and $\ln(10^{10}A_{s})$, as illustrated in the $\ln(10^{10}A_{s}){-}f_{c}$ plane shown in the right panel of Figure˜6. When fixing the parameter $f_{c}$ to its expected value, referred to the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model (yellow), we successfully recover a constraining power comparable to that of the clean case. The estimated mean value of $\ln(10^{10}A_{s})$ inferred by the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model shows a slightly larger deviation ($\sim 0.85\sigma$), likely due to the combined effects of redshift uncertainty in the slitless-like errors and the approximation applied in the correction model (see the $R(k)$ comparison in Figure˜3).
This test is informative for slitless spectroscopy survey like Euclid, which aim to constrain cosmological parameters via galaxy power spectra. As we have shown that this hypothetical redshift errors can affect our cosmological inference, an accurate estimation of the catastrophics rate $f_{c}$ and the implementation of the catastrophics-corrected model, either with $f_{c}$ fixed or varied under a narrow Gaussian prior, will be essential to ensure unbiased constraints.
@@
Figure 6: Posterior distributions for Full-Modeling cosmological parameters. Left panel: Constraints using the baseline EFT model to fit clean mocks (dark blue), mocks contaminated by ELG-like catastrophics ($f_{c}=1\%$; purple), and slitless-like errors ($\sigma_{v}=85.7\rm{km}\cdot s^{-1}$ with $f_{c}=5\%$; red). Right panel: Comparison between clean results with slitless-like errors using the extended EFT model when freeing $f_{c}$ ($\texttt{EFT}+{\rm{free}}\,f_{c}$; orange) and with $f_{c}$ fixed to its expected value ($\texttt{EFT}+{\rm{fixed}}\,f_{c}$; yellow). Contours show 68% and 95% confidence intervals with black dash lines indicating the true values. Results are derived from fits to the mean power spectrum monopole and quadrupole of 500 mocks using the covariance matrix rescaled to $V_{25}$. The most notable change is in $\ln(10^{10}A_{s})$ estimate, for which the constraints are summarized in Table˜2.
@@
Theory EFT baseline $\texttt{EFT}+{\rm{free}}\,f_{c}$ $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ Mocks clean dv-ELGcatas dv-slitless dv-slitless dv-slitless $\ln(10^{10}A_{s})$ $3.125\pm 0.084$ $3.095\pm 0.081$ $2.941\pm 0.085$ $3.13\pm 0.14$ $3.054\pm 0.080$
@@
Table 2: Constraints on $\ln(10^{10}A_{s})$ for clean and contaminated mocks analyzed by different theoretical models. Table shows the mean value with 68% confidence intervals. In EFT baseline fit, the ELG-like catastrophics (dv-ELGcatas) and sliltless-like errors (dv-slitless) show deviations of $0.36\sigma$ and $2.20\sigma$, respectively, compared to the clean case. By multiplying the factor $(1-f_{c})^{2}$ to the galaxy power spectrum, the extended model $f_{c}$, $\texttt{EFT}+{\rm{free}}\,f_{c}$ recovers an unbiased value of $\ln(10^{10}A_{s})$ but increases its error bar by 60% due to parameter degeneracies. In contract, the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model restore the posterior without the loss of constraining power.
@@ 5.3 Beyond $\Lambda\rm{CDM}$ cosmologies
Recent spectroscopic surveys such as DESI have extended full-shape power spectrum analyses beyond the standard $\Lambda\rm{CDM}$ framework to test cosmological models with evolving dark energy ($w_{0}w_{a}\rm{CDM}$) and massive neutrinos ($\nu\Lambda\rm{CDM}$). Assuming a flat Universe and combining the full-shape and BAO analysis with external datasets, the DESI DR2 analysis finds a mild preference for a dynamical dark energy model with $w_{0}>-1$ and $w_{a}<0$  [5] and places a tight constraint on the total neutrino mass, $\sum m_{\nu}<0.0642\,\rm{eV}\ (95\%CL)$ [65]. This section investigates how redshift errors affect the measurements of dark energy equation of state $w_{0}$-$w_{a}$ and $\sum m_{\nu}$ through the full-shape analysis. In the following, our results show fits to the mean power spectrum of 500 mocks with the covariance rescaled to volume $V_{25}$.
@@ 5.3.1 Evolving dark energy
Figure 7: Parameter constraints on $w_{0}w_{a}\rm{CDM}$ model for clean and contaminated mocks. The horizontal error bars indicate mean and the 68% credible levels of the corresponding posterior. The shaded region show the $1\sigma$ constraints from the clean mocks. The open circles represent the MAP values. While redshift errors do not introduce bias in the measurement of dark energy parameters $w_{0}$, $w_{a}$, they potentially lead to stronger projection effects (deviation between MAP and mean of the posterior), particularly in cases with larger redshift uncertainty.
We assess the influence of redshift errors on the measurement of evolving dark energy using the Chevallier-Polarski-Linder (CPL) parametrization $w(a)=w_{0}+w_{a}(1-a)$ [66]. The cosmological fits are performed using the Quijote fiducial mocks described in Section˜3, which assume a $\Lambda\rm{CDM}$ model with $w_{0}=-1$ and $w_{a}=0$. The resulting parameter constraints are shown as the horizontal lines in Figure˜7. From the comparisons, we find that redshift errors do not introduce significant bias in the measurement of dark energy properties. However, we know that using full-shape analyses alone to constraint $w_{0}w_{a}\rm{CDM}$ model suffer from projection effects [9]. To study this aspect, we use iminuit [67] to estimate the maximum a posteriori (MAP) value, starting from the maximum log-posterior points found in MCMC chains. In Figure˜7, we find offsets between the mean of the marginalised posteriors (solid circles) and MAP values (open circles) for parameters $h$, $w_{0}$ and $w_{a}$, with the offsets being most pronounced for mocks contaminated by QSO-like Lorentzian smearing. Since the redshift errors impact the clustering pattern similarly with the behavior of certain parameters (cosmological or nuissance), their presence could complicate the parameter degeneracies and thus amplify projection effects. Such impact should be investigated in more detail for redshift surveys aiming to constrain dark energy by full-shape analysis.
@@ 5.3.2 Massive neutrinos
Figure 8: 1D marginalized posterior distributions of sum of neutrino mass, $\sum m_{\nu}$, for uncontaminated mock (solid darkblue line) and contaminated (dashed lines) mock. Left panel show cases with LRG-like (purple), QSO-like (Gaussian, light purple; Lorentzian, orange) redshift uncertainties. Right panel use contaminated mocks with ELG-like catastrophic ($f_{c}=1\%$, blue) and slitless-like errors ($\Delta v_{\rm{smear}}=85.7\rm{km}\cdot s^{-1}$ with $f_{c}=5\%$, green). The shaded region indicates the 95% confidence level from the clean mocks, while the vertical dotted line marks the true value $\sum m_{\nu}=0.1\,\rm{eV}$.
Catalogs clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless $\sum m_{\nu}$ $0.80\,\rm{eV} <$$1.07\,\rm{eV} <$$1.08\,\rm{eV} <$$1.03\,\rm{eV} <$$0.88\,\rm{eV} <$$0.83\,\rm{eV}$
Table 3: Comparison of 95% confidence upper bounds on the sum of neutrino masses, $\sum m_{\nu}$. Results are obtained using Full-Modeling analyses with $\sum m_{\nu}$ varied and fitting power spectrum monopole and quadrupole of 500 mocks with covariance rescaled to $V_{25}$. Both clean and contaminated mock catalogs are considered with $\sum m_{\nu}=0.1,\mathrm{eV}$. We find that redshift uncertainties degrade the neutrino mass constraint by $\sim 30\%$, whereas ELG-like redshift catastrophics and slitless-like errors show negligible effect.
For the $\nu\Lambda\rm{CDM}$ case, we use the Quijote Mnu_pMnu\_p catalogs with $\sum m_{\nu}=0.1\,\rm{eV}$ to build our clean and contaminated mock catalogs. During the fitting, we allow $\sum m_{\nu}$ to vary with a flat prior of $\mathcal{U}(0.0,5.0)\,\rm{eV}$, while keeping other settings the same as the baseline. The marginalized posteriors of $\sum m_{\nu}$ for all mocks analyses are displayed in Figure˜8 ana the upper limits of $\sum m_{\nu}$ at 95% confidence level are presented in Table˜3. We first note that using full-shape analysis alone to constrain the $\sum m_{\nu}$ has slow convergence and large uncertainties, particularly for the low neutrino mass case. Nevertheless, we find that redshift uncertainty introduces extended tails in the $\sum m_{\nu}$ posterior distribution, leading to a weaker constraint on the neutrino mass. Compared to the clean case ($0.80\,\rm{eV}$), the 95% upper limits increase by $\sim 28$ to 35% when redshift uncertainty is included. This degradation occurs because both massive neutrinos and redshift uncertainty have similar suppression on small-scale power spectrum. Massive neutrinos do so through their thermal velocities and free-streaming, while redshift uncertainty smear the clustering structure along the line of sight. This degeneracy complicates the convergence and leads to weaker constraints on neutrino mass. In contrast, we find no noticeable bias in the $\sum m_{\nu}$ measurements due to ELG-like catastrophic failures or slitles-like redshift errors.
@@ 6 Conclusions
In this paper, we investigate the impact of spectroscopic redshift errors on the cosmological measurements. Section˜2 introduces two main types of redshift errors, redshift uncertainty and redshift catastrophic failures, which can affect redshift measurements in both ground-based (fiber-fed) and space-based (sltiless) spectroscopic surveys such as eBOSS, DESI and Euclid [35, 21, 17]. While redshift uncertainty refers to small errors present in all redshift measurements, redshift catastrophic failures correspond to large deviations from the true values in a small fraction of the measurements. Both types of error distort the observed clustering pattern and can lead to potential biases in the inferred cosmological parameters.
@@
The effect of redshift errors can be quantified and modeled using repeat redshift observations. Based on this, in Section˜3 we use the Quijote simulation suite to constructed our contaminated mock catalogs. For redshift uncertainty, we consider three cases: LRG-like Gaussian, QSO-like Gaussian and QSO-like Lorentzian smearings, with the velocity dispersions informed by DESI EDR studies [26]. For redshift catastrophics, we use a log-normal Gaussian distribution [21] with contamiantion rate $f_{c}=1\%$ to model ELG-like catastrophics and $f_{c}=5\%$ for hypothetical slitless-like errors. The slitless-like errors also combines LRG-like Gaussian smearing to simulate the observing condition expected in slitless spectroscopic survey such as Euclid. The following Section˜4 describes the methodologies to perform the full-shape power spectrum analysis. We use clustering models based on the effective field theory (EFT) of LSS and implement both ShapeFit and Full-Modeling approaches to extract cosmological information, and we fit the power spectrum monopole and quadrupole for $k\in[0.02,0.20]\,h\,\rm{Mpc}^{-1}$. The analysis is targeted on DESI- and Euclid-like survey with effective volume $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$. Our key findings are summarized below:
@@
Redshift uncertainty introduces diffusive effect in the clustering, leading to a scale-dependent suppression to the observed power spectrum. This smearing effect can be modeled by a Gaussian damping term [20]. We find that the EFT cournterterms, parameterized by $\alpha_{2}$ and $\alpha_{4}$, also capture this smearing effect, resulting in compressed and cosmological parameter estimates difference $5\%$ compared to the uncontaminated case. In contrast, redshift catastrophics cause an overall reduction on the power spectrum amplitude that approximately scales as $(1-f_{c})^{2}$. While this effect is negligible (difference $6\%$) for ELG-like catastrophics, it significantly biases the estimation of amplitude-related parameters for slitless-like errors, namely $df$ in ShapeFit and $\ln(10^{10}A_{s})$ in the Full-Modeling. In that case, we observe differences of $6\%{-}16\%$, corresponding to a $\sim 2.2\sigma$ shift relative to the clean case. To model the effect of redshift catastrophics, we use the correction factor $(1-f_{c})^{2}$ applied to the power spectrum in Eq.˜4.6. When freeing $f_{c}$, the extended model mitigates the bias in $\ln(10^{10}A_{s})$ to below $0.1\sigma$, but leads to $60\%$ degradation in its precision due to the strong degeneracy between $f_{c}$ and $\ln(10^{10}A_{s})$. Fixing $f_{c}$ to its expected value in the fit restores the constraining power, with only a modest $0.85\sigma$ deviation. This highlights the importance of accurately estimating the catastrophic failure rate $f_{c}$ and applying the appropriate correction model to ensure unbiased cosmological results in slitless spectroscopic surveys.
@@
Future work should focus on refining the model of redshift errors, exploring higher-order statistics to complement the power spectrum analyses, using light-cone mocks and more observational data to validate and extend these findings. In particular, we plan to apply our framework to DESI DR2 analysis, which provides a large sample of repeat observations well-suited for characterizing redshift errors. Using this dataset, we aim to jointly model redshift uncertainty and redshift catastrophics for each tracer population, across a wide range of redshift, enabling a more realistic and tracer-specific treatment of redshift systematics for cosmological inference.

### v4

@@ 5.2.2 Catastrophics-correction for slitless survey
Figure 8: Posterior distributions for Full-Modeling parameters. Left panel: Constraints using the baseline EFT model to fit clean mocks (dark blue), mocks contaminated by ELG-like catastrophics ($f_{c}=1\%$; blue), and slitless-like errors ($\sigma_{v}=85.7\rm{km}\cdot s^{-1}$ with $f_{c}=5\%$; green). Right panel: Comparison between clean results with slitless-like errors using the extended EFT model when freeing $f_{c}$ ($\texttt{EFT}+{\rm{free}}\,f_{c}$; orange) and with $f_{c}$ fixed to its expected value ($\texttt{EFT}+{\rm{fixed}}\,f_{c}$; yellow). Contours show 68% and 95% confidence intervals with black dash lines indicating the true values. Results are derived from fits to the mean power spectrum monopole and quadrupole of 500 mocks using the covariance matrix rescaled to $V_{25}$. The marginalized posteriors of $\ln(10^{10}A_{s})$ and $b_{1}$ are shifted by slitless-like errors, but using the corrected model with $f_{c}$ help restore the estimated values.
@@
Theory EFT baseline $\texttt{EFT}+{\rm{free}}\,f_{c}$ $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ Mocks clean dv-ELGcatas dv-slitless dv-slitless dv-slitless $\ln(10^{10}A_{s})$ $3.125\pm 0.084$ $3.095\pm 0.081$ $2.941\pm 0.085$ $3.13\pm 0.14$ $3.042\pm 0.079$
@@
Table 2: Constraints on $\ln(10^{10}A_{s})$ for clean and contaminated mocks fitted by different theoretical models. Table shows the mean values with 68% confidence intervals. In EFT baseline fit, the mean values from ELG-like catastrophics and sliltless-like errors show deviations of $0.4\sigma$ and $2.2\sigma$, respectively, compared to the clean case. By multiplying the factor $(1-f_{c})^{2}$ to the galaxy power spectrum, the extended model $f_{c}$, $\texttt{EFT}+{\rm{free}}\,f_{c}$ recovers an unbiased value of $\ln(10^{10}A_{s})$ but increases its error bar by 67% due to parameter degeneracies. Fixing the $f_{c}$ to its expected value, the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model restores the constraining power with a modest bias of $1.0\sigma$.
In this test, we present Full-Modeling fitting results based on the mean power spectrum calculated from all 500 mocks, with the covariance matrix rescaled to volume $V_{25}$. By averaging over 500 mocks, we minimize the impact of cosmic variance, so any observed deviations in parameter estimates can be more confidently attributed to redshift errors. The left panel of Figure˜8 shows the posterior distributions of parameters using the baseline EFT fit for three contamination cases: clean, ELG-like catastrophics ($f_{c}=1\%$), and slitless-like errors (i.e., $\sigma_{v}=85.7\rm{km}\cdot s^{-1}$ and $f_{c}=5\%$). The corresponding mean values of $\ln(10^{10}A_{s})$ with 68% confidence intervals are summarized in Table˜2. Comparing the $\ln(10^{10}A_{s})$ constraint from clean case, we find a small shift of $0.4\sigma$ by ELG-like catastrophics and a significant one of $2.2\sigma$ by slitless-like errors. These results are consistent with our earlier findings in section˜5.2.1. The linear bias parameter $b_{1}$ is also known to be related to the overall amplitude of galaxy power spectrum. In Figure˜8, we find a shift of $2.7\sigma$ in its mean value for the slitless-like errors case using baseline EFT. We note that $b_{1}$ is highly degenerate with $f_{c}$ and $\ln(10^{10}A_{s})$, but the shift in $b_{1}$ is insufficient to compensate for the bias induced by slitless-like errors on $\ln(10^{10}A_{s})$.
To incorporate the effect of catastrophics in the clustering model, we extend the baseline EFT model by adding an additional parameter, catastrophic rate $f_{c}$, and multiplying the factor $(1-f_{c})^{2}$ to the galaxy power spectrum as described in Eq.˜4.7. When freeing $f_{c}$ parameter during sampling with a flat prior $f_{c}\in[0,1]$, we refer to the $\texttt{EFT}+{\rm{free}}\,f_{c}$ model. The posterior distributions, including the variation of $f_{c}$, are shown in the right panel of Figure˜8, with the corresponding constraint on $\ln(10^{10}A_{s})$ reported in Table˜2. The $\texttt{EFT}+{\rm{free}}\,f_{c}$ model succeed in recovering an unbiased mean value of $\ln(10^{10}A_{s})$ for slitless-like errors, with a discrepancy below $0.1\sigma$ compared to the clean case. However, this improvement comes at the cost of loss in constraining power: we measure a $\sim 67\%$ degradation in the $\ln(10^{10}A_{s})$ constraint. This loss is due to the strong degeneracy between $f_{c}$ and $\ln(10^{10}A_{s})$, as illustrated in the Figure˜8 $\ln(10^{10}A_{s}){-}f_{c}$ plane. When fixing the parameter $f_{c}$ to its expected value, referred to the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model, we successfully recover the constraining power but find a slightly larger deviation ($\sim 1.0\sigma$) in the estimated mean value. This deviation is possibly due to the combined effects of redshift uncertainty in the slitless-like errors and the approximation applied in the correction model.
This test is informative for slitless spectroscopy survey like Euclid, which aims to constrain cosmological parameters via galaxy power spectra. As we have shown that this hypothetical redshift errors can bias cosmological inference, an accurate estimation of the catastrophics rate and the implementation of the catastrophics-corrected model, either with $f_{c}$ fixed or varied under a narrow Gaussian prior, are necessary to get unbiased constraints. Since our slitless-like errors scenario is optimistic and slitless survey may have other additional complex systematics, further investigations on modeling redshift errors are still required.
@@ 5.3 Beyond $\Lambda\rm{CDM}$ cosmologies
Recent spectroscopic surveys such as DESI have extended full-shape power spectrum analysis beyond the standard $\Lambda\rm{CDM}$ framework to test cosmological models with evolving dark energy ($w_{0}w_{a}\rm{CDM}$) and massive neutrinos. Assuming a flat Universe and combining the full-shape and BAO analyses with external datasets, the DESI DR2 finds a mild preference for a dynamical dark energy model with $w_{0}>-1$ and $w_{a}<0$  [desi_collaboration_desi_2025-4] and places a tight constraint on the total neutrino mass, $\sum m_{\nu}<0.0642\,\rm{eV}\ (95\%CL)$ [elbers_constraints_2025]. This section investigates how redshift errors affect the measurements of dark energy equation of state $w_{0}$-$w_{a}$ and $\sum m_{\nu}$ through the full-shape analysis. In the following, we use the baseline EFT fitting setup and present results obtained by fitting the mean power spectrum of 500 mocks, with the covariance matrix rescaled to the effective volume $V_{25}$.
@@ 5.3.1 Evolving dark energy
Figure 9: Parameter constraints on $w_{0}w_{a}\rm{CDM}$ model for clean and contaminated mocks. The horizontal error bars indicate mean and the 68% credible levels of the corresponding posterior. The shaded region show the $1\sigma$ constraints from the clean mocks. While redshift errors do not introduce bias in the measurement of dark energy properties $w_{0}$, $w_{a}$, they potentially lead to stronger projection effects, which are deviations between the MAP values (open circles) with the mean of the posterior, particularly with QSO-like smearing.
We assess the influence of redshift errors on the measurement of evolving dark energy using the Chevallier-Polarski-Linder (CPL) parametrization $w(a)=w_{0}+w_{a}(1-a)$ [chevallier_accelerating_2001, linder_exploring_2003]. The cosmological fits are performed using the Quijote fiducial mocks described in section˜3, which assume a $\Lambda\rm{CDM}$ model with $w_{0}=-1$ and $w_{a}=0$. The resulting parameter constraints are shown as the horizontal lines in Figure˜9. From the comparisons, we find that redshift errors do not introduce significant bias in the measurement of dark energy properties. However, we know that using full-shape analysis alone to constraint $w_{0}w_{a}\rm{CDM}$ model suffer from projection effects [desi_collaboration_desi_2025-1]. To study this aspect, we use iminuit [james_minuit_1975] to estimate the maximum a posteriori (MAP) value, starting from the maximum log-posterior points found in MCMC chains. In Figure˜9, we find offsets between the mean of the marginalised posteriors (solid circles) and MAP values (open circles) for parameters $h$, $w_{0}$ and $w_{a}$, with the offsets being most pronounced for mocks contaminated by QSO-like smearing. Since the redshift errors impact the clustering pattern similarly with the behavior of certain parameters (cosmological or nuisance), their presence could complicate the parameter degeneracies and thus amplify projection effects. Such impact should be investigated in more detail for redshift surveys aiming to constrain dark energy by full-shape analysis.
@@ 5.3.2 Massive neutrinos
Catalogs clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless $\sum m_{\nu}$ $0.78\,\rm{eV} <$$0.69\,\rm{eV} <$$1.40\,\rm{eV} <$$0.94\,\rm{eV} <$$0.84\,\rm{eV} <$$0.85\,\rm{eV}$
Table 3: Comparison of 95% confidence upper bounds on the sum of neutrino masses, $\sum m_{\nu}$. Results are obtained using Full-Modeling analysis with $\sum m_{\nu}$ varied and fitting power spectrum monopole and quadrupole of 500 mocks with covariance rescaled to $V_{25}$. Both clean and contaminated mock catalogs are considered with $\sum m_{\nu}=0.1\,\mathrm{eV}$. We find that redshift uncertainties degrade the neutrino mass constraint up to $\sim 80\%$ in QSO-like smearing cases, whereas ELG-like redshift catastrophics and slitless-like errors show negligible effect.
For the massive neutrino case, we use the Quijote Mnu_pMnu\_p catalogs with $\sum m_{\nu}=0.1\,\rm{eV}$ to build our clean and contaminated mock catalogs. During the fitting, we allow $\sum m_{\nu}$ to vary with a flat prior of $\mathcal{U}(0.0,5.0)\,\rm{eV}$ in the $\Lambda\rm{CDM}$ framework. The upper limits of $\sum m_{\nu}$ at 95% confidence level for all mocks analyses are presented in Table˜3. We first note that using full-shape analysis alone to constrain the $\sum m_{\nu}$ has slow convergence and large uncertainties, particularly for the low neutrino mass case. Nevertheless, we find that redshift uncertainty introduces extended tails in the $\sum m_{\nu}$ posterior distribution, leading to a weaker constraint on the neutrino mass. Compared to the clean case ($0.78\,\rm{eV}$), the 95% upper limits increase by up to $\sim 80\%$ when QSO Gaussian smearing is included. This degradation occurs because both massive neutrinos and redshift uncertainty have similar suppression on small-scale power spectrum. Massive neutrinos do so through their thermal velocities and free-streaming, while redshift uncertainty smears the clustering structure along the line of sight. This degeneracy complicates the convergence and leads to weaker constraints on neutrino mass. In contrast, we find no noticeable bias in the $\sum m_{\nu}$ measurements due to ELG-like catastrophic failures or slitles-like redshift errors.
@@ 6 Conclusions
In this paper, we investigate the impact of spectroscopic redshift errors on cosmological measurements. section˜2 introduces the redshift errors, which can affect redshift measurements in both ground-based (fiber-fed) and space-based (slitless) spectroscopic surveys [ross_completed_2020, yu_elg_2025, euclid_collaboration_euclid_2025-1]. There are two main types of redshift errors. Redshift uncertainty refer to small errors present in all redshift measurements, and redshift catastrophic failures correspond to large deviations from the true values in a small fraction of the measurements. Both type of errors distort the observed clustering pattern and can lead to biases in the inferred cosmological results.
@@
The effect of redshift errors can be quantified and modeled using repeat observations. In section˜3 we use the Quijote simulation suite to construct our contaminated mock catalogs. For redshift uncertainty, we consider Gaussian or Lorentzian profiles for LRG-like and QSO-like smearing [zarrouk_clustering_2018, yu_desi_2023]. For redshift catastrophics, we use a log-normal distribution [yu_elg_2025] with contamination rate $f_{c}=1\%$ to model ELG-like catastrophics. The slitless-like errors assume a hypothetical scenario combining $f_{c}=5\%$ catastrophics and LRG-like uncertainty to simulate the observing condition in slitless survey. The following section˜4 describes the methodologies to perform the full-shape power spectrum analysis. In our baseline fitting, we use clustering models based on the effective field theory (EFT) of LSS, without including the damping term or the multiplicative factor. We implement both ShapeFit and Full-Modeling fitting approach to extract cosmological information. The analysis is targeted on DESI- and Euclid-like surveys with effective volume $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$. Our key findings are summarized below:
@@
Redshift uncertainty introduces diffusive effect in the clustering, leading to a scale-dependent suppression to the observed power spectrum. The EFT cournterterms, parameterized by $\alpha_{2}$ and $\alpha_{4}$, capture this smearing effect, resulting in compressed and cosmological parameter estimates difference $5\%$ compared to the uncontaminated case. In contrast, redshift catastrophics cause an overall reduction on the power spectrum amplitude that approximately scales as $(1-f_{c})^{2}$. While this effect is negligible (difference $6\%$) for ELG-like catastrophics, it significantly biases the estimation of amplitude-related parameters for slitless-like errors, namely $df$ in ShapeFit and $\ln(10^{10}A_{s})$ in the Full-Modeling. In that case, we observe shifts of $6\%{-}16\%$, corresponding to a $\sim 2.2\sigma$ deviation relative to the clean case. To model this effect, we use the correction factor $(1-f_{c})^{2}$ applied to the power spectrum in Eq.˜4.7. When freeing $f_{c}$, the extended model mitigates the bias in $\ln(10^{10}A_{s})$ to below $0.1\sigma$, but leads to $67\%$ degradation in its precision due to the strong degeneracy between $f_{c}$ and $\ln(10^{10}A_{s})$. Fixing $f_{c}$ to its expected value in the fit restores the constraining power with a modest $1.0\sigma$ deviation. This highlights the importance of accurately estimating the catastrophic failure rate $f_{c}$ and applying the appropriate correction model to ensure unbiased cosmological results in slitless spectroscopic surveys.
@@
For future work, we plan to apply our framework to the DESI DR2 analysis. The DESI DR2 dataset provides a large sample of repeat observations well-suited for characterizing redshift errors. Using this dataset, we aim to jointly model redshift uncertainty and catastrophic failures for each tracer population across a wide range of redshift. We also plan to conduct more detailed investigations of redshift errors, including lightcone effects, redshift evolution, impact on all various clustering statistics and so on.